# How Do Recommendation Models Amplify Popularity Bias? An Analysis from the Spectral Perspective

### v3

@@ Abstract
Recommendation Systems (RS) are often plagued by popularity bias. When training a recommendation model on a typically long-tailed dataset, the model tends to not only inherit this bias but often exacerbate it, resulting in over-representation of popular items in the recommendation lists. This study conducts comprehensive empirical and theoretical analyses to expose the root causes of this phenomenon, yielding two core insights: 1) Item popularity is memorized in the principal spectrum of the score matrix predicted by the recommendation model; 2) The dimension collapse phenomenon amplifies the relative prominence of the principal spectrum, thereby intensifying the popularity bias.
@@
Building on these insights, we propose a novel debiasing strategy that leverages a spectral norm regularizer to penalize the magnitude of the principal singular value. We have developed an efficient algorithm to expedite the calculation of the spectral norm by exploiting the spectral property of the score matrix. Extensive experiments across seven real-world datasets and three testing paradigms have been conducted to validate the superiority of the proposed method.
1 Introduction
@@
Recommender Systems (RS), with their capability to offer personalized suggestions, have found applications across various domains [18, 44, 72]. Nevertheless, their effectiveness in personalization is significantly compromised by popularity bias [12]. This bias emerges when recommendation data showcases a long-tailed distribution of item interaction frequencies. Subsequently, recommendation models trained on such data tend to inherit and even amplify this bias, leading to an overwhelming presence of popular items in recommendation results [56, 68, 74]. This notorious effect not only undermines the accuracy and fairness of recommendation [5, 4], but also exacerbates the Matthew Effect and the filter bubble through the user-system feedback loop [37, 23, 22].
@@
Given the detrimental impact of popularity bias amplification, a thorough understanding of its root causes is crucial. Although some recent studies have endeavored to elucidate this, their investigations exhibit significant limitations: 1) Some researchers [56, 68, 55] have investigated popularity bias amplification through causal graphs. However, they merely postulate causal relations between item popularity and model predictions without deeply exploring the underlying mechanisms behind the relations. Moreover, their analyses depend on hypothesized causal graphs, which may be flawed due to the widespread presence of unmeasured confounders [21, 67]. 2) Other studies [73, 69, 11, 65] revealed that graph neural networks (GNNs) can exacerbate popularity bias. However, these analyses are limited to GNNs as applied in specific graph-based recommendation models, rather than the mechanisms of generic recommendation models.
@@
2) Amplification Effect. The phenomenon known as dimension collapse augments the relatively prominence of the principal spectrum that captures item popularity, leading to bias amplification (Figure 1(b)). We reveal that dimension collapse is pervasive in RS due to two primary reasons: i) The deliberate low-rank setting of user/item embeddings, employed either to conserve memory or to counteract overfitting, amplifies the impact of the principal spectrum; ii) The inherent training dynamics of gradient-based optimization prioritize the learning of the principal dimension, while the singular values of other dimensions are easily underestimated. Our further theoretical and empirical analyses establish the relationship between dimension collapse and popularity bias — larger principal singular values compared to other singular values lead to more popular items on the recommendations.
@@
Our analysis not only explains the underlying mechanisms of bias amplification but also paves the way for the development of an innovative strategy to counteract this effect. Recognizing that the essence of this amplification lies in the undue contribution of the principal spectrum, we introduce a spectral norm regularizer [60] aimed at directly restraining the magnitude of the principal singular value. However, the direct computation of the spectral norm necessitates exhaustive processing of a large score matrix and numerous iterative procedures [60, 52], inducing significant computational costs. To address this challenge, we further develop an accelerated strategy by leveraging the intrinsic spectrum properties of the score matrix and matrix transformation techniques. Consequently, our method effectively mitigates popularity bias while imposing limited computational overhead. It is worth noting that our method is model-agnostic and can be easily integrated into various recommendation models. We conduct extensive experiments on seven real-world datasets with three representative backbones, validating the superiority of our proposed method over existing debiasing methods.
Figure 1: Illustration of two core insights.
2 Preliminaries
@@
Task Formulation. This work mainly focus on the collaborative filtering (CF) [58], a widely-used recommendation scenario. Consider a RS with a user set $\mathcal{U}$ and an item set $\mathcal{I}$. Let $n$ and $m$ denote the total number of users and items. Historical interactions can be expressed by a matrix $\mathbf{Y}\in\{0,1\}^{n\times m}$, where the element $y_{ui}$ indicates if user $u$ has interacted with item $i$ (e.g., click). For convenience, we define the number of interactions of an item as $r_{i}=\sum_{u\in\mathcal{U}}{y_{ui}}$, and collect $r_{i}$ over all items as a popularity vector $\mathbf{r}$. RS targets to suggest items to users based on their potential interests.
@@
Recommendation Models. Embedding-based models are widely utilized in RS [58]. Such models convert user/item attributes (e.g., IDs) into $d$-dimensional representations ($\mathbf{u}_{u},\mathbf{v}_{i}$), and make predictions using the embedding similarity [58]. Given that the inner product is a conventional similarity metric due to its efficiency in retrieval and superior performance [57, 61, 34], this work also focuses on the inner product for analysis. Specifically, the model’s predicted scores can be formulated as $\hat{y}_{ui}=\mu(\mathbf{u}_{u}^{\top}\mathbf{v}_{i})$, where $\mu(.)$ denotes an activation function like Sigmoid. $\hat{y}_{ui}$ represents a user’s preference for an item, which is then used for ranking to generate recommendations. For clarity of presentation, we also employ matrix notation. Let matrices $\hat{\mathbf{Y}}$, ${\mathbf{U}}$, ${\mathbf{V}}$ represent scores over all user-item combinations, embeddings over all users and items, respectively. Model predictions can be succinctly expressed as ${\bf{\hat{Y}}}=\mu({\bf{U}}{{\bf{V}}^{\top}})$.
Objective Functions. Common choices of loss functions for training a recommendation model include point-wise loss such as BCE and MSE [45], and pair-wise loss like BPR [46]. It is worth noting that BPR can be reconceptualized as a specialized pointwise loss if we construct hyper-item spaces and consider a pair of items as a unique hyper-item (please refer to Appendix A.1.1 for more details). As such, for convenience, this work focuses on point-wise loss for analysis. But we also discuss how our proposed debiased method adapts to BPR loss (Appendix A.1.2) and validate its effectiveness through experiments (Section 5).
@@
Popularity Bias Amplification. Items’ interaction frequency in recommendation data often follows a long-tailed distribution [8, 17, 51]. For instance, in a typical Douban dataset, a mere 20% of the most popular items account for 86.3% of all interactions. When models are trained on such skewed data, they tend to absorb and amplify this bias, frequently over-prioritizing popular items in their recommendations. For example, in the Douban dataset using the MF model, 20% of the most popular items occupy over 99.7% of the recommendation slots, while a mere 0.6% of the most popular items occupy more than 63% (cf. Appendix B.1 more examples). This notorious effect significantly impacts the recommendation accuracy and fairness, even potentially posing detrimental effects on the entire ecosystem of RS[12]. Thus, understanding the underlying mechanisms behind this effect is crucial.
3 Understanding Popularity Bias Amplification
3.1 Popularity Bias Memorization Effect
@@
3.1.1 Empirical Study. To discern how recommendation models memorize item popularity, we designed the following experiment: 1) We well trained three representative recommendation models, MF [38], LightGCN [27] and XSimGCL [62], on three real-world datasets (cf. Section 5 for experimental details); 2) We then performed SVD decomposition on the predicted score matrix, $\hat{\mathbf{Y}}=\mathbf{P}\mathbf{\Sigma}\mathbf{Q}^{\top}=\sum_{1\leq k\leq L}{{\sigma_{k}}{{\bf{p}}_{k}}{\bf{q}}_{k}^{\top}}$ where $L=min(n,m)$ and $\sigma_{1}\geq\sigma_{2}\geq...\geq\sigma_{L}$. We further computed cosine similarity between the right principal singular vector $\mathbf{q}_{1}$ and the item popularity $\mathbf{r}$. The outcomes are showcased in Table LABEL:cos_sim. From these experiments, we draw an impressive observation:
@@
Table 1: The cosine similarity between the principal singular vector ($\mathbf{q}_{1}$) and the item popularity ($\mathbf{r}$) under different backbones and loss functions.
@@
Here $r_{g}$ signifies the popularity of the $g$-th most popular item, and $\alpha$ is a shape parameter indicating the distribution’s slope. Power-law, as a typical long-tailed distribution, is prevalent across various natural and man-made phenomena [16]. Recent studies assert that item popularity in RS also aligns with this ubiquitous principle [8, 17, 51]. Then we have the following important theorem:
@@
$\cos({\bf{r}},{{\bf{q}}_{1}})\geq\frac{{\sigma_{1}^{2}}}{{{r_{\max}}\sqrt{\zeta(2\alpha)}}}\sqrt{1-\frac{{{r_{\max}}(\zeta(\alpha)-1)}}{{\sigma_{1}^{2}}}}$ (1)
@@
$\cos({\bf{r}},{{\bf{q}}_{1}})\geq\sqrt{\frac{{2-\zeta(\alpha)}}{{\zeta(2\alpha)}}}$ (2)
@@
Proof can be found in Appendix A.2. Notably, as the long-tailed nature of item popularity intensifies (i.e., $\alpha\to\infty$ suggesting $\zeta(\alpha)\to 1$), the right side of Eq. (2) converges to one, implying a near-perfect alignment between $\bf{r}$ and $\bf q_{1}$. Even when the data isn’t markedly skewed and has a considerable $\zeta(\alpha)$, we typically observe $\sigma_{1}^{2}$ to vastly exceed $r_{\max}$, e.g., ${\rm{5.6}\times 10^{5}}$ vs. ${\rm{4.6}\times 10^{3}}$ in the dataset Movielens (with more examples presented in Appendix B.2). Thus, from Eq. (1), a high similarity between $\bf{r}$ and $\bf q_{1}$ emerges. This theorem provides theoretical validation for our observation 1.
(a) Exploring dimensions.
@@
Figure 2: Illustration of how dimension collapse impacts popularity bias in Movielens: (a)-(b) the proportion of popular items in recommendations and the ratio of the largest singular value ($\sigma_{1}^{2}/\sum_{1\leq k\leq L}{\sigma_{k}^{2}}$) with varying embedding dimensions and training epochs, respectively; (c) how singular values evolves during training.
3.2 Popularity Bias Amplification Effect
Earlier discussions illuminate that the principal spectrum memorizes item popularity. In this subsection, we reveal the phenomenon of dimension collapse in recommendation systems (RS), which amplifies the effect of the principal spectrum, leading to popularity bias amplification.
3.2.1 Empirical Study. The occurrence of dimension collapse in RS is largely attributable to two factors: 1) explicit low-rank configuration of user/item embeddings [39, 27], and 2) intrinsic training dynamics associated with gradient-based optimization [15, 48, 6]. Here, we present experiments to validate these points and examine their impacts on popularity bias.

### v6

@@ Abstract.
Recommendation Systems (RS) are often plagued by popularity bias. When training a recommendation model on a typically long-tailed dataset, the model tends to not only inherit this bias but often exacerbate it, resulting in over-representation of popular items in the recommendation lists. This study conducts comprehensive empirical and theoretical analyses to expose the root causes of this phenomenon, yielding two core insights: 1) Item popularity is memorized in the principal spectrum of the score matrix predicted by the recommendation model; 2) The dimension reduction phenomenon amplifies the relative prominence of the principal spectrum, thereby intensifying the popularity bias.
@@
Building on these insights, we propose a novel debiasing strategy that leverages a spectral norm regularizer to penalize the magnitude of the principal singular value. We have developed an efficient algorithm to expedite the calculation of the spectral norm by exploiting the spectral property of the score matrix. Extensive experiments across seven real-world datasets and three testing paradigms have been conducted to validate the superiority of the proposed method. The code is available at https://github.com/LIN452/ReSN/.
Recommender System; Popularity Bias
journalyear: 2025copyright: acmlicensedconference: Proceedings of the Eighteenth ACM International Conference on Web Search and Data Mining; March 10–14, 2025; Hannover, Germany.booktitle: Proceedings of the Eighteenth ACM International Conference on Web Search and Data Mining (WSDM ’25), March 10–14, 2025, Hannover, Germanyisbn: 979-8-4007-1329-3/25/03doi: 10.1145/3701551.3703579ccs: Information systems Recommender systems
1. Introduction
@@
Recommender Systems (RS), with their capability to offer personalized suggestions, have found applications across various domains (Covington et al., 2016; Qiu et al., 2018; Zhou et al., 2018). Collaborative filtering (CF), a widely-used technique within RS, learns user preference from historical interactions. However, their effectiveness in personalization is significantly compromised by popularity bias (Chen et al., 2023a). This bias emerges when user interaction data showcases a long-tailed distribution of item interaction frequencies. Subsequently, recommendation models trained on such data tend to inherit and even amplify this bias, leading to an overwhelming presence of popular items in recommendation results (Wei et al., 2021; Zhang et al., 2021; Zhu et al., 2021).
This notorious effect not only undermines the accuracy and fairness of recommendation (Abdollahpouri et al., 2020, 2019), but also exacerbates the Matthew Effect and the filter bubble through the user-system feedback loop (Mansoury et al., 2020; Gao et al., 2023b, a).
@@
Given the detrimental impact of popularity bias amplification, a thorough understanding of its root causes is crucial. Although some recent studies have endeavored to elucidate this, their investigations exhibit significant limitations: 1) Some researchers (Wei et al., 2021; Zhang et al., 2021; Wang et al., 2021) have investigated popularity bias amplification through causal graphs. However, they merely postulate causal relations between item popularity and model predictions without deeply exploring the underlying mechanisms behind the relations. Moreover, their analyses depend on hypothesized causal graphs, which may be flawed due to the widespread presence of unmeasured confounders (Ding et al., 2022; Zhang et al., 2018). 2) Other studies (Zhou et al., 2023; Zhang et al., 2023b, 2024; Chen et al., 2024b, a) have revealed graph neural network (GNNs) can exacerbate popularity bias. However, these analyses primarily focus to GNNs rather than the mechanisms of generic recommendation models.
@@
2) Amplification Effect. The phenomenon known as dimension reduction augments the relatively prominence of the principal spectrum that captures item popularity, leading to bias amplification (Figure 1(b)). We reveal that dimension reduction is pervasive in RS due to two primary reasons: i) The deliberate low-rank setting of user/item embeddings, employed either to conserve memory or to counteract overfitting, amplifies the impact of the principal spectrum; ii) The inherent training dynamics of gradient-based optimization prioritize the learning of the principal dimension, while the singular values of other dimensions are easily underestimated. Our further theoretical and empirical analyses establish the relationship between dimension reduction and popularity bias — larger principal singular values compared to other singular values lead to more popular items on the recommendations.
Figure 1. Illustration of two core insights.
@@
Our analysis not only explains the underlying mechanisms of bias amplification but also paves the way for the development of an innovative strategy to counteract this effect. Recognizing that the essence of this amplification lies in the undue contribution of the principal spectrum, we introduce a spectral norm regularizer (Yoshida and Miyato, 2017) aimed at directly restraining the magnitude of the principal singular value. However, the direct computation of the spectral norm necessitates exhaustive processing of a large score matrix and numerous iterative procedures (Yoshida and Miyato, 2017; Sun et al., 2021), inducing significant computational costs. To address this challenge, we further develop an accelerated strategy by leveraging the intrinsic spectrum properties of the score matrix and matrix transformation techniques. Consequently, our method effectively mitigates popularity bias while imposing limited computational overhead.
In summary, our contributions are:
  • Conducting comprehensive analyses to unravel the mechanisms behind popularity bias amplification in recommendation — item popularity is encoded within the principal singular vector, and its impact is exaggerated due to the dimension reduction phenomenon.
  • Proposing an efficient method for mitigating the bias amplification through the regulation of the principal singular value.
  • Performing extensive experiments across seven real-world datasets under three different testing scenarios, demonstrating the superiority of our method in reducing bias and enhancing recommendation quality.
2. Preliminaries
In this section, we present the background of the recommendation system and popularity bias amplification.
@@ 2. Preliminaries
Task Formulation. This work mainly focus on the collaborative filtering (CF) (Wu et al., 2022), a widely-used recommendation scenario. Consider a RS with a user set $\mathcal{U}$ and an item set $\mathcal{I}$. Let $n$ and $m$ denote the total number of users and items. Historical interactions can be expressed by a matrix $\mathbf{Y}\in\{0,1\}^{n\times m}$, where the element $y_{ui}$ indicates if user $u$ has interacted with item $i$ (e.g., click). For convenience, we define the number of interactions of an item as $r_{i}=\sum_{u\in\mathcal{U}}{y_{ui}}$, and collect $r_{i}$ over all items as a popularity vector $\mathbf{r}$. RS targets to suggest items to users based on their potential interests.
@@
Recommendation Models. Embedding-based models are widely utilized in RS (Wu et al., 2022). Such models convert user/item attributes (e.g., IDs) into $d$-dimensional representations ($\mathbf{u}_{u},\mathbf{v}_{i}$), and make predictions using the embedding similarity (Wu et al., 2022). Given that the inner product is a conventional similarity metric due to its efficiency in retrieval and superior performance (Wu et al., 2021; Yu et al., 2022; Lin et al., 2022), this work also focuses on the inner product for analysis. Specifically, the model’s predicted scores can be formulated as $\hat{y}_{ui}=\mu(\mathbf{u}_{u}^{\top}\mathbf{v}_{i})$, where $\mu(.)$ denotes an activation function like Sigmoid. $\hat{y}_{ui}$ represents a user’s preference for an item, which is then used for ranking to generate recommendations. For clarity of presentation, we also employ matrix notation. Let matrices $\hat{\mathbf{Y}}$, ${\mathbf{U}}$, ${\mathbf{V}}$ represent scores over all user-item combinations, embeddings over all users and items, respectively. Model predictions can be succinctly expressed as ${\mathbf{\hat{Y}}}=\mu({\mathbf{U}}{{\mathbf{V}}^{\top}})$.
Objective Functions. Common choices of loss functions for training a recommendation model include point-wise loss such as BCE and MSE (Rendle, 2022), and pair-wise loss like BPR (Rendle et al., 2009). It is worth noting that BPR can be reconceptualized as a specialized pointwise loss.Concretely, BPR loss is expressed as:
${\mathcal{L}_{BPR}}=-\sum\limits_{u\in\mathcal{U}}{\sum\limits_{i\in\mathcal{I},{y_{ui}}=1}{\sum\limits_{j\in\mathcal{I},{y_{uj}}=0}\log({\mu({\mathbf{u}}_{u}^{\top}{{\mathbf{v}}_{i}}-{\mathbf{u}}_{u}^{\top}{{\mathbf{v}}_{j}})})}}\vspace{-4pt}$
If construct a hyper-item space denoted as $\mathcal{I}^{\prime}=\mathcal{I}\times\mathcal{I}$ derived from item pairs, and define the embeddings of hyper-items as ${\mathbf{v}^{\prime}_{ij}}={{\mathbf{v}}_{i}}-{{\mathbf{v}}_{j}}$ and assign new observed interactions to the combinations of users and hyper-items, i.e., ${y^{\prime}_{u,ij}}=1$ for $y_{ui}=1\&y_{uj}=0$ and ${y^{\prime}_{u,ij}}=0$ for $y_{ui}=0\&y_{uj}=1$. BPR can be re-written as:
$(\displaystyle{\mathcal{L}_{BPR}}=-\frac{1}{2}\sum\limits_{u\in\mathcal{U}}\Big{(}$$\sum_{\begin{subarray}{c}(i,j)\in\mathcal{I^{\prime}}\\ y^{\prime}_{u,ij}=1\end{subarray}}\log(\mu({\mathbf{u}}_{u}^{\top}{\mathbf{v}^{\prime}_{ij}}))+\sum_{\begin{subarray}{c}(i,j)\in\mathcal{I^{\prime}}\\ y^{\prime}_{u,ij}=0\end{subarray}}\log(\mu(-{\mathbf{u}}_{u}^{\top}{\mathbf{v}^{\prime}_{ij}}))\Big{)}$
where BPR can be reframed as a specific point-wise loss under the hyper-items space $\mathcal{I^{\prime}}$. Therefore, for convenience, our analyses mainly focus on point-wise loss. But we will also discuss why our proposed debiased method is suitable for BPR (cf. Section 4.2) and validate its effectiveness in experiments (cf. Section 5).
@@
Popularity Bias Amplification. Items’ interaction frequency in recommendation data often follows a long-tailed distribution (Borges and Stefanidis, 2020; Çoba et al., 2017; Steck, 2011). For instance, in a typical Douban dataset, a mere 20% of the most popular items account for 86.3% of all interactions. When models are trained on such skewed data, they tend to absorb and amplify this bias, frequently over-prioritizing popular items in their recommendations. For example, in the Douban dataset using the MF model, 20% of the most popular items occupy over 99.7% of the recommendation slots, while a mere 0.6% of the most popular items occupy more than 63% (cf. Appendix B.1 more examples). This notorious effect significantly impacts the recommendation accuracy and fairness, even potentially posing detrimental effects on the entire ecosystem of RS(Chen et al., 2023a). Thus, understanding the underlying mechanisms behind this effect is crucial.
3. Understanding Popularity Bias Amplification
In this section, we conduct thorough analyses to answer:
1) How do recommendation models memorize the item popularity?
2) Why do recommendation models amplify popularity bias?
3.1. Popularity Bias Memorization Effect
@@
3.1.1 Empirical Study. To discern how recommendation models memorize item popularity, we designed the following experiment: 1) We well trained three representative recommendation models, MF (Marlin and Zemel, 2009), LightGCN (He et al., 2020) and XSimGCL (Yu et al., 2023), on three real-world datasets (cf. Section 5 for experimental details); 2) We then performed SVD decomposition on the predicted score matrix, $\hat{\mathbf{Y}}=\mathbf{P}\mathbf{\Sigma}\mathbf{Q}^{\top}=\sum_{1\leq k\leq L}{{\sigma_{k}}{{\mathbf{p}}_{k}}{\mathbf{q}}_{k}^{\top}}$ where $L=min(n,m)$ and $\sigma_{1}\geq\sigma_{2}\geq...\geq\sigma_{L}$. We further computed cosine similarity between the right principal singular vector $\mathbf{q}_{1}$ and the item popularity $\mathbf{r}$. The outcomes are showcased in Table LABEL:cos_sim. From these experiments, we draw an impressive observation:
Observation 1.
@@
Table 1. The cosine similarity between the principal singular vector ($\mathbf{q}_{1}$) and the item popularity ($\mathbf{r}$) under different backbones and loss functions.
@@
Here $r_{g}$ signifies the popularity of the $g$-th most popular item, and $\alpha$ is a shape parameter indicating the distribution’s slope. Power-law, as a typical long-tailed distribution, is prevalent across various natural and man-made phenomena (Clauset et al., 2009). Recent studies assert that item popularity in RS also aligns with this ubiquitous principle (Borges and Stefanidis, 2020; Çoba et al., 2017; Steck, 2011). Then we have the following important theorem:
@@
(1) $\cos({\mathbf{r}},{{\mathbf{q}}_{1}})\geq\frac{{\sigma_{1}^{2}}}{{{r_{\max}}\sqrt{\zeta(2\alpha)}}}\sqrt{1-\frac{{{r_{\max}}(\zeta(\alpha)-1)}}{{\sigma_{1}^{2}}}}$
@@
(2) $\cos({\mathbf{r}},{{\mathbf{q}}_{1}})\geq\sqrt{\frac{{2-\zeta(\alpha)}}{{\zeta(2\alpha)}}}$
@@
Proof can be found in Appendix A.1. Notably, as the long-tailed nature of item popularity intensifies (i.e., $\alpha\to\infty$ suggesting $\zeta(\alpha)\to 1$), the right side of Eq. (2) converges to one, implying a near-perfect alignment between $\mathbf{r}$ and $\mathbf{q}_{1}$. Even when the data isn’t markedly skewed and has a considerable $\zeta(\alpha)$, we typically observe $\sigma_{1}^{2}$ to vastly exceed $r_{\max}$, e.g., ${\rm{5.6}\times 10^{5}}$ vs. ${\rm{4.6}\times 10^{3}}$ in the dataset Movielens (with more examples presented in Appendix B.2). Thus, from Eq. (1), a high similarity between $\mathbf{r}$ and $\mathbf{q}_{1}$ emerges. This theorem provides theoretical validation for our observation 1.
(a) Exploring different dimensions.
@@
Figure 2. Illustration of how dimension reduction impacts popularity bias in Movielens: (a)-(b) the proportion of popular items in recommendations and the ratio of the largest singular value ($\sigma_{1}^{2}/\sum_{1\leq k\leq L}{\sigma_{k}^{2}}$) with varying embedding dimensions and training epochs, respectively; (c) how singular values evolves during training.
3.2. Popularity Bias Amplification Effect
Earlier discussions illuminate that the principal spectrum memorizes item popularity. In this subsection, we reveal the phenomenon of dimension reduction in RS, which amplifies the effect of the principal spectrum, leading to popularity bias amplification.
3.2.1 Empirical Study. The occurrence of dimension reduction in RS is largely attributable to two factors: 1) explicit low-rank configuration of user/item embeddings (Mnih and Salakhutdinov, 2007; He et al., 2020), and 2) intrinsic training dynamics associated with gradient-based optimization (Chou et al., 2023; Saxe et al., 2019; Arora et al., 2019). Here, we present experiments to validate these points and examine their impacts on popularity bias.